Caspian Sea Drinks is considering the production of a diet drink. The expansion of the plant and the purchase of the equipment necessary to produce the diet drink will cost $28.00 million. The plant and equipment will be depreciated over 10 years to a book value of $3.00 million, and sold for that amount in year 10. Net working capital will increase by $1.31 million at the beginning of the project and will be recovered at the end. The new diet drink will produce revenues of $8.85 million per year and cost $1.91 million per year over the 10-year life of the project. Marketing estimates 15.00% of the buyers of the diet drink will be people who will switch from the regular drink. The marginal tax rate is 33.00%. The WACC is 12.00%. Find the IRR (internal rate of return).

Submit

Answer format: Percentage Round to: 4 decimal places (Example: 9.2434%, % sign required. Will accept decimal format rounded to 6 decimal places (ex: 0.092434))
Answer #1

IRR=11.26%

Workings

Incremental profit = (revenue-cost)*(1-Switching customers)

= (8850000-1910000)*(1-15%)

= 5899000

Incremental profit after tax = 5899000*(1-33%) = 3952330

Year Initial cost Tax shield Salvage Net workings capital Incremental profit after tax Net Cash flows
0 -28000000 -1310000 -29310000
1 825000 3952330 4777330
2 825000 3952330 4777330
3 825000 3952330 4777330
4 825000 3952330 4777330
5 825000 3952330 4777330
6 825000 3952330 4777330
7 825000 3952330 4777330
8 825000 3952330 4777330
9 825000 3952330 4777330
10 825000 3000000 1310000 3952330 9087330

Book1- Excel AutoSave Off Sign in Share View O Tell me what you want to do File Review Help Comments Home Insert Draw Page La